OpenStudy (harrisonshark):

Simplify. a - {4b - [a - (5b - c) + 3c - (a - b - 9c)]}

OpenStudy (harrisonshark):

please help

OpenStudy (harrisonshark):

options: -a - 8b + 13c a - 8b + 13c a - 2b + 13c -a - 2b + 13c

OpenStudy (anonymous):

a - {4b - [a - (5b - c) + 3c - (a - b - 9c)]} =a - {4b - [a - 5b + c + 3c - a + b + 9c]} =a-{4b-[-4b+13c]} =a-{4b+4b-13c} =a-8b+13c answer is option A

OpenStudy (harrisonshark):

u mean b?

OpenStudy (anonymous):

sorrey.answer is option B...not A...